OpenStudy (anonymous):

Write an equation and solve. At the local warehouse where Billy buys supplies for the concession stand, a case of energy drinks costs twice as much as a case of water and a case of chips together. If a case of energy drinks costs $18 and a case of water costs $5, what is the cost of a case of chips?

OpenStudy (tyteen4a03):

Let the cost of a case of chips be x. The equation you're looking for is 2(5+x) = 18 The 5 stands for the cost for a case of water, and 18 stands for the cost for a case of energy drinks.

OpenStudy (anonymous):

let the case of energy drinks = x let the case of water = y case of chips together = z x = (y+z)*2 substitute the values

OpenStudy (anonymous):

I understand all that, but what is the answer?

OpenStudy (tyteen4a03):

@HorseCrazyGirlForever The answer is given at my post.

OpenStudy (anonymous):

18=(5+z)2 z=4

OpenStudy (anonymous):

18?